# Module-conditioned distribution of quantum circuits

### v1

@@ 1 Introduction
This paper primarily considers homogeneous networks with $k\geq 3$ modules. In this setting, distributions using a hypergraph partitioning formulation and heuristic solvers proved to be the most effective [3, 22]. However, this approach jointly determines the module allocation and the implementation of non-local gates based on that allocation, even though the former cannot logically depend on the latter. As our main contribution, we observe that exact integer programming formulations [23, 24, 25] can be derived if the module allocation function is fixed, demonstrating stability for various types of circuits. First, we present formulations for $k\geq 4$ modules and then show how it simplifies for $k=3$ modules. As a result, our method can be combined with any DQC solver by ignoring its distribution steps and adopting only the module allocation it returns. Our approach further reduces ebit costs for various types of circuits, with significant improvements observed in some structured circuits. This highlights the importance of our post-processing for achieving robust results.
@@ 2.2 Cat-entanglement & cat-disentanglement
between two modules $A$ and $B$, which is a maximally entangled bipartite state. In the context of DQC, such states are referred to as ebits. Figure 1 shows how to implement a non-local controlled unitary gate with a single control qubit $|c\rangle$ and a single target qubit $|t\rangle$ at the cost of a single ebit. The steps before (after) the controlled unitary gate constitute a cat-entanglement (cat-disentanglement) process. The former creates a linked copy of $|c\rangle$ in module $B$, and the latter effectively reverses this process. After cat-disentanglement, the operation $|ct\rangle\mapsto CP(\theta)|ct\rangle$ has been successfully implemented.
@@ 2.4 Hypergraph partitioning formulation
Figure 2 shows an example circuit distributed by this method. The coverage of the gate $\alpha$ corresponds to 3 in Definition 2, and the coverages of the gates $\beta$ and $\gamma$ corresponds to 1 or 2 in Definition 2.
@@ 3.1 BIP formulation of MS-GC for $k\geq 4$
Proof. We show that it suffices to relax the constraint
@@
Table 1: Truth table for (3.1) & (6).
0 0 0 T T
0 0 1 F F
0 1 0 T T
0 1 1 F F
1 0 0 T T
1 0 1 F F
1 1 0 T T
1 1 1 T T
The only combination that violates (3) while making (6) true is
and
i.e., $\hat{x}^{(\textbf{m})}$ is also an optimal solution of (3.1), which satisfies (3). $\blacksquare$
@@ 3.2 BIP formulation of MS-GC for $k=3$
Proof. We prove by contradiction. Let $P=\{p_{1},p_{2},p_{3}\}$ and $\pi(q_{i})=p_{1}$. Suppose $m=(q_{i},p_{2},t)\notin M^{(1)}$ for some $m\in M$ is strictly necessary. By the definition of $M^{(1)}$, there is no gate ($(\{q_{i},q_{j}\},t^{\prime})$ such that $\pi(q_{j})=p_{2}$ and $f(i,t^{\prime})=t$. This means that the non-local binary gates that $m$ can cover (by itself or in a pair with another migration) must be in the form of ($(\{q_{i},q_{j}\},t^{\prime})$, where $\pi(q_{j})=p_{3}$ and $f(i,t^{\prime})=t$ (and in fact, $m$ cannot cover one of these gates by itself because it migrates $q_{i}$ in $p_{1}$ to $p_{2}$ while the gate acts on a qubit in $p_{3}$; it must be paired with a migration that sends $q_{j}$ to $p_{2}$). Since we assumed that $m$ is strictly necessary, there must exist at least one such gate. Meanwhile, notice that those gates are all covered by the single migration $m^{\prime}=(q_{i},p_{3},t)\in M^{(1)}$. Therefore, replacing $m$ with $m^{\prime}$ does not increase the number of selected migrations without introducing any uncovered gates, which implies that any migration $m\notin M^{(1)}$ is not strictly necessary; each of them can be replaced by some $m^{\prime}\in M^{(1)}$. $\blacksquare$
Figure 3 provides an example circuit to visualize Lemma 1, as well as a circuit with $k=4$ to which the lemma does not apply.
@@
Proof. By the definition of $A$, it is clear that $A_{i,:}x\geq 2$ if and only if a non-empty subset of $\left\{M^{(1,B_{i})}_{1},M^{(1,B_{i})}_{2},M^{(2,B_{i})}_{1}\right\}$ is selected, where $A_{i,:}$ denotes the $i$-th row of $A$. $\blacksquare$
@@ 3.3 Examples and analysis
Figure 4 shows an example circuit distributed by the hypergraph partitioning formulation introduced in Section 2.4 and the BIP formulation described in this section. Even for these small circuits, the hypergraph partitioner begins to find suboptimal solutions, and the BIP post-processing helps reduce the ebit cost.
@@ 3.4 Partitioning QFT
The basic structure of QFT before distribution is shown in Figure 5. It consists of the Hadamard gates $H$ and $CP$ gates
@@
Proof. We provide a simple constructive proof. The QFT circuit for $n=km$ qubits can be fully described by the sets
@@
because the circuit includes only one unary gate per qubit and one binary gate per pair of qubits. That is, $t_{q_{i}}$ and $t_{\{q_{i},q_{j}\}}$ are uniquely defined. Assume that the qubits and gates are arranged in the conventional order as shown in Figure 5 and let $\pi$ be a balanced module allocator function, i.e.,
@@
covers all gates in $B$ with ebit cost $k(k-1)m/2$, which achieves its lower bound (9). $\blacksquare$
Note that $\pi^{*}$ in Lemma 2 is not the only optimal module allocation for MS-HC (see Figure 6). However, for $(k,m)=(3,2)$, it is very likely that $\pi^{*}$ is the only optimal module allocation for MS-GC (see Table 2). We also ran exhaustive experiments to find the optimal allocations among all balanced allocations for $(k,m)=(4,2)$ and $(k,m)=(3,3)$ (see Figure 7). In both cases, the canonical partitions $\pi^{*}$ were optimal for MS-GC (assuming that the BIP solver found optimal solutions for these small examples). That is, distributing a QFT circuit by solving the BIP associated with $\pi^{*}$ is expected to yield the lowest ebit cost we can hope for. Indeed, we observe that this heuristic proves effective (see Section 4.2.1).
@@
4 Experiments
4.1 Random circuits
4.1.1 $CZ$ fraction circuits
Given parameters $(n,d,p)$, an $n$-qubit $CZ$ fraction circuit is generated as described in Algorithm 3 [1]. Figure 8 shows an example $CZ$ fraction circuit. In all plots, ‘HP’ stands for the hypergraph partitioning solution. Ebit costs for distributing random $CZ$ fraction circuits are shown in Figure 10.
4.1.2 Quantum volume circuits
Given parameters $(n,d)$, an $n$-qubit quantum volume circuit is generated as described in Algorithm 4 [29]. Figure 9 (right) shows an example quantum volume circuit. Ebit costs for distributing random quantum volume circuits are shown in Figure 11.
@@ 4.2.1 QFT circuits
Ebit costs for distributing QFT circuits are shown in Figure 12. We remark that replacing $\pi$ (found by a hypergraph partitioner) with $\pi^{*}$ defined in (10) improved the ebit cost in every single run of the BIP solver. In this regard, we conjecture that for the distribution of QFT circuits, $\pi^{*}$ is not only the optimal module allocation function for MS-HC but also for MS-GC.
@@ 4.2.2 DraperQFTAdder
where the subscripts denote qubit indices (see APPENDIX C). The right-hand side appears many times in DraperQFTAdder circuits (Figure 13). Since SWAP gates require binary gates to implement, we replace the right-hand side of (12) with the left-hand side. Ebit costs for distributing DraperQFTAdder circuits are shown in Figure 14.
@@ 4.2.3 RGQFTMultiplier
The RGQFTMultiplier is a quantum circuit that leverages QFT to perform multiplication [27]. They store the product of two $n^{\prime}$-bit inputs out-of-place. By default, the output register has $2n^{\prime}$ qubits, resulting in a total of $4n^{\prime}$ qubits (Figure 15).
Ebit costs for distributing RGQFTMultiplier circuits are shown in Figure 16. Surprisingly, the hypergraph partitioner alone returns extremely inefficient distributions in many cases; post-processing with the BIP solver reduces the ebit costs significantly.
@@ 4.3.1 AND circuits
The AND circuit implements the logical AND operation on a number of qubits, i.e., it is a multi-controlled $X$ gate (Figure 17a). Ebit costs for distributing AND circuits are shown in Figure 18.
@@ 4.3.2 InnerProduct circuits
Ebit costs for distributing InnerProduct circuits are shown in Figure 19.
[…]
$\pi$ ebit cost $\pi$ ebit cost
112233 4 112323 5
112332 5 121233 5
121323 6 121332 6
122133 5 123123 6
123132 6 122313 6
123213 6 123312 6
122331 5 123231 6
123321 6 -- --
Figure 8: An example $7$-qubit $3$-layer $CZ$ fraction circuit with $p=4/7$.
[…]
Figure 11: BIP improvements observed in quantum volume circuits. The reduction in ebit cost is relatively marginal for these circuits. This is because as mentioned in [22], rewriting a given quantum volume circuit such that all non-unary gates are CP gates makes the fraction of binary gates quite low. Similar to the results for circuits with low $CZ\_\text{percent}$ shown in Figure 10, little improvement is expected.
[…]
Figure 12: BIP and/or $\pi^{*}$ improvements observed in QFT circuits.
[…]
Figure 14: BIP improvements observed in DraperQFTAdder circuits.
Figure 16: BIP improvements observed in RGQFTMultiplier circuits.
[…]
Figure 18: BIP improvements observed in AND circuits.
[…]

### v2

@@ Abstract
keywords:
Distributed quantum computing, Integer programming
Acknowledgments
This work is in part supported by the National Research Foundation of Korea (NRF, RS-2024-00451435 (20%), RS-2024-00413957 (40%)), Institute of Information & Communications Technology Planning & Evaluation (IITP, 2021-0-01059 (40%)), grant funded by the Ministry of Science and ICT (MSIT), Institute of New Media and Communications (INMAC), and the Brain Korea 21 FOUR program of the Education and Research Program for Future ICT Pioneers.
@@ 1 Introduction
This paper primarily considers homogeneous networks with $k\geq 3$ modules. In this setting, distributions using a hypergraph partitioning formulation and heuristic solvers proved to be the most effective [32, 1]. However, this approach jointly determines the module allocation and the implementation of non-local gates based on that allocation, even though the former cannot logically depend on the latter. As our main contribution, we observe that exact integer programming formulations [26, 20, 31] yield improved ebit costs if the module allocation function is fixed, demonstrating stability for various types of circuits. First, we present formulations for $k\geq 4$ modules and then show how it simplifies for $k=3$ modules. As a result, our method can be combined with any DQC solver by ignoring its distribution steps and adopting only the module allocation it returns. Our approach further reduces ebit costs for various types of circuits, with significant improvements observed in some structured circuits. This highlights the importance of our post-processing for achieving robust results.
@@ 2.2 Cat-entanglement & cat-disentanglement
between two modules $A$ and $B$, which is a maximally entangled bipartite state. In the context of DQC, such states are referred to as ebits. Fig. 1 shows how to implement a non-local controlled unitary gate with a single control qubit $|c\rangle$ and a single target qubit $|t\rangle$ at the cost of a single ebit. The steps before (after) the controlled unitary gate constitute a cat-entanglement (cat-disentanglement) process. The former creates a linked copy of $|c\rangle$ in module $B$, and the latter effectively reverses this process. After cat-disentanglement, the operation $|ct\rangle\mapsto CP(\theta)|ct\rangle$ has been successfully implemented.
@@ 2.4 Hypergraph partitioning formulation
Fig. 2 shows an example circuit distributed by this method. The coverage of the gate $\alpha$ corresponds to 3 in Definition 2, and the coverages of the gates $\beta$ and $\gamma$ corresponds to 1 or 2 in Definition 2.
@@ 3.1 BIP formulation of MS-GC for $k\geq 4$
Proof.
We show that it suffices to relax the constraint
@@
Table 1: Truth table for (3.1) & (6).
0 0 0 T T
0 0 1 F F
0 1 0 T T
0 1 1 F F
1 0 0 T T
1 0 1 F F
1 1 0 T T
1 1 1 T T
The only combination that violates (3) while making (6) true is
and
i.e., $\hat{x}^{(\textbf{m})}$ is also an optimal solution of (3.1), which satisfies (3). ∎
@@ 3.2 BIP formulation of MS-GC for $k=3$
Proof.
We prove by contradiction. Let $P=\{p_{1},p_{2},p_{3}\}$ and $\pi(q_{i})=p_{1}$. Suppose $m=(q_{i},p_{2},t)\notin M^{(1)}$ for some $m\in M$ is strictly necessary. By the definition of $M^{(1)}$, there is no gate ($(\{q_{i},q_{j}\},t^{\prime})$ such that $\pi(q_{j})=p_{2}$ and $f(i,t^{\prime})=t$. This means that the non-local binary gates that $m$ can cover (by itself or in a pair with another migration) must be in the form of ($(\{q_{i},q_{j}\},t^{\prime})$, where $\pi(q_{j})=p_{3}$ and $f(i,t^{\prime})=t$ (and in fact, $m$ cannot cover one of these gates by itself because it migrates $q_{i}$ in $p_{1}$ to $p_{2}$ while the gate acts on a qubit in $p_{3}$; it must be paired with a migration that sends $q_{j}$ to $p_{2}$). Since we assumed that $m$ is strictly necessary, there must exist at least one such gate. Meanwhile, notice that those gates are all covered by the single migration $m^{\prime}=(q_{i},p_{3},t)\in M^{(1)}$. Therefore, replacing $m$ with $m^{\prime}$ does not increase the number of selected migrations without introducing any uncovered gates, which implies that any migration $m\notin M^{(1)}$ is not strictly necessary; each of them can be replaced by some $m^{\prime}\in M^{(1)}$. ∎
Fig. 3 provides an example circuit to visualize Lemma 1, as well as a circuit with $k=4$ to which the lemma does not apply.
@@
Proof.
By the definition of $A$, it is clear that $A_{i,:}x\geq 2$ if and only if a non-empty subset of $\left\{M^{(1,B_{i})}_{1},M^{(1,B_{i})}_{2},M^{(2,B_{i})}_{1}\right\}$ is selected, where $A_{i,:}$ denotes the $i$-th row of $A$. ∎
@@ 3.3 Examples and analysis
Fig. 4 shows an example circuit distributed by the hypergraph partitioning formulation introduced in Section 2.4 and the BIP formulation described in this section. Even for these small circuits, the hypergraph partitioner begins to find suboptimal solutions, and the BIP post-processing helps reduce the ebit cost.
@@ 3.4 Partitioning QFT
The basic structure of QFT before distribution is shown in Fig. 5. It consists of the Hadamard gates $H$ and $CP$ gates
@@
Proof.
We provide a simple constructive proof. The QFT circuit for $n=km$ qubits can be fully described by the sets
@@
because the circuit includes only one unary gate per qubit and one binary gate per pair of qubits. That is, $t_{q_{i}}$ and $t_{\{q_{i},q_{j}\}}$ are uniquely defined. Assume that the qubits and gates are arranged in the conventional order as shown in Fig. 5 and let $\pi$ be a balanced module allocator function, i.e.,
@@
covers all gates in $B$ with ebit cost $k(k-1)m/2$, which achieves its lower bound (9). ∎
$\pi$ ebit cost $\pi$ ebit cost
112233 4 112323 5
112332 5 121233 5
121323 6 121332 6
122133 5 123123 6
123132 6 122313 6
123213 6 123312 6
122331 5 123231 6
123321 6 -- --
Note that $\pi^{*}$ in Lemma 2 is not the only optimal module allocation for MS-HC (see Fig. 6). However, for $(k,m)=(3,2)$, it is very likely that $\pi^{*}$ is the only optimal module allocation for MS-GC (see Table 2). We also ran exhaustive experiments to find the optimal allocations among all balanced allocations for $(k,m)=(4,2)$ and $(k,m)=(3,3)$ (see Fig. 7). In both cases, the canonical partitions $\pi^{*}$ were optimal for MS-GC (assuming that the BIP solver found optimal solutions for these small examples). That is, distributing a QFT circuit by solving the BIP associated with $\pi^{*}$ is expected to yield the lowest ebit cost we can hope for. Indeed, we observe that this heuristic proves effective (see Section 4.2.1).
@@ 4.1.1 $CZ$ fraction circuits
Figure 8: An example $7$-qubit $3$-layer $CZ$ fraction circuit with $p=4/7$.
Given parameters $(n,d,p)$, an $n$-qubit $CZ$ fraction circuit is generated as described in Algorithm 3 [17]. Fig. 8 shows an example $CZ$ fraction circuit. In all plots, ‘HP’ stands for the hypergraph partitioning solution. Ebit costs for distributing random $CZ$ fraction circuits are shown in Fig. 9.
4.1.2 Quantum volume circuits
@@
Figure 11: BIP improvements observed in quantum volume circuits. The reduction in ebit cost is relatively marginal for these circuits. This is because as mentioned in [1], rewriting a given quantum volume circuit such that all non-unary gates are CP gates makes the fraction of binary gates quite low. Similar to the results for circuits with low $CZ\_\text{percent}$ shown in Fig. 9, little improvement is expected.
Given parameters $(n,d)$, an $n$-qubit quantum volume circuit is generated as described in Algorithm 4 [11]. Fig. 10 (right) shows an example quantum volume circuit. Ebit costs for distributing random quantum volume circuits are shown in Fig. 11.
@@ 4.2.1 QFT circuits
Ebit costs for distributing QFT circuits are shown in Fig. 12. We remark that replacing $\pi$ (found by a hypergraph partitioner) with $\pi^{*}$ defined in (10) improved the ebit cost in every single run of the BIP solver. In this regard, we conjecture that for the distribution of QFT circuits, $\pi^{*}$ is not only the optimal module allocation function for MS-HC but also for MS-GC.
Figure 12: BIP and/or $\pi^{*}$ improvements observed in QFT circuits.
@@ 4.2.2 DraperQFTAdder
Figure 14: BIP improvements observed in DraperQFTAdder circuits.
where the subscripts denote qubit indices (see APPENDIX C). The right-hand side appears many times in DraperQFTAdder circuits (Fig. 13). Since SWAP gates require binary gates to implement, we replace the right-hand side of (12) with the left-hand side. Ebit costs for distributing DraperQFTAdder circuits are shown in Fig. 14.
@@ 4.2.3 RGQFTMultiplier
The RGQFTMultiplier is a quantum circuit that leverages QFT to perform multiplication [27]. They store the product of two $n^{\prime}$-bit inputs out-of-place. By default, the output register has $2n^{\prime}$ qubits, resulting in a total of $4n^{\prime}$ qubits (Fig. 15).
Figure 16: BIP improvements observed in RGQFTMultiplier circuits.
Ebit costs for distributing RGQFTMultiplier circuits are shown in Fig. 16. Surprisingly, the hypergraph partitioner alone returns extremely inefficient distributions in many cases; post-processing with the BIP solver reduces the ebit costs significantly.
@@ 4.3.1 AND circuits
Figure 18: BIP improvements observed in AND circuits.
The AND circuit implements the logical AND operation on a number of qubits, i.e., it is a multi-controlled $X$ gate (Fig. 17a). Ebit costs for distributing AND circuits are shown in Fig. 18.
@@ 4.3.2 InnerProduct circuits
Ebit costs for distributing InnerProduct circuits are shown in Fig. 19.